If the rate on one-year T-Bills currently is 2 percent, what is the repayment probability for each of the following two securities?

Assume that if the loan is defaulted, no payments are expected. What is the market-determined probability of default for each security?

a) One-year AA-rated bond yielding 7.5 percent.

b) One-year BB-rated bond yielding 10 percent.
Answer #1

1.
(1+2%)=(1-pd)*(1+7.5%)
=>pd=1-(1+2%)/(1+7.5%)
=>pd=5.12%

2.
(1+2%)=(1-pd)*(1+10%)
=>pd=1-(1+2%)/(1+10%)
=>pd=7.27%